OpenStudy (anonymous):

A rocket travels vertically at a speed of 1200 km/hr. The rocket is tracked through a telescope by an observer located 11 km from the launching pad. Find the rate at which the angle between the telescope and the ground is increasing 3 min after lift-off. (Round your answer to two decimal places.)

OpenStudy (perl):

my draw button is not working

OpenStudy (anonymous):

Where to go from there?

OpenStudy (perl):



OpenStudy (perl):

\( \Large { \tan\theta = \frac{y}{11}\\ ~ \\ \\ \therefore implicit ~ derivative~ with~ respect ~ to ~ time \\ \sec^2 \theta \cdot \frac{d\theta}{dt} = \frac{1}{11} \cdot \frac{dy}{dt} } \)

OpenStudy (perl):

we want to find \( \Large \frac{d\theta}{dt} \) , when \( \Large \frac{dy}{dt}=1200 \) and t = 3/60 hour

OpenStudy (anonymous):

What is d theta/dt?